OpenStudy (anonymous):

Simplify 9+3[5+(6)^2] please explain!

OpenStudy (souvik):

6^2=?

OpenStudy (anonymous):

36

OpenStudy (souvik):

well...now[5+6^2]=?

OpenStudy (anonymous):

41

OpenStudy (souvik):

good...now 3*[5+6^2]=?

OpenStudy (anonymous):

where did you get the 3?.. D0 you mean 9+3?..

OpenStudy (souvik):

look at the question...3 is multiplied with [5+6^2]

OpenStudy (anonymous):

ohh my bad, 123?

OpenStudy (souvik):

yes...now 9+3[5+(6)^2] =?

OpenStudy (anonymous):

132 thanks (:

OpenStudy (souvik):

yep! right...you are welcome...:)